OpenStudy (superhelp101):

Which point lies on the graph of the following piecewise function? y = x2 if x less than or equal to 0 y = x + 1 if x > 0 (0, 1) (-1, 0) (-2, 4) (1, 1)

OpenStudy (superhelp101):

@thomaster

OpenStudy (superhelp101):

@aum can you help me?

OpenStudy (superhelp101):

@dan @lacrosseplayer22

OpenStudy (superhelp101):

@mathstudent55

OpenStudy (anonymous):

Evaluate the function for the \(x\) coordinate, then see if it matches up with the listed \(y\) coordinate.

OpenStudy (superhelp101):

@SithsAndGiggles could you explain to me?

OpenStudy (anonymous):

For instance, check \((1,1)\). \(x=1\), so and you plug it into \(f(x)\) and you get \(f(x)=y=1\), then this point is on the graph. \(x=1\) is greater than 0, so you use the linear part, \(f(x)=x+1\). \[f(1)=1+1=2\] When \(x=1\), \(y=2\), so \((1,1)\) is not on the graph.

OpenStudy (superhelp101):

oh could I just draw the tables

OpenStudy (anonymous):

What do you mean by tables?

OpenStudy (superhelp101):

I learned a way with table where you use numbers

OpenStudy (superhelp101):

lol that's not helpful, but let you use that then we can see what you did and compare methods. Just a sec :)

OpenStudy (anonymous):

It sounds like the same method in principle, you plug in numbers and arrange them in a table.

OpenStudy (superhelp101):

oh I got 1,1 was that right

OpenStudy (anonymous):

No, I eliminated that one. \(1>0\), so you use \(f(x)=x+1\).

OpenStudy (superhelp101):

umm. hmm

OpenStudy (superhelp101):

I plugged in different numbers into each function and got that both functions are the same at 1, 1

OpenStudy (superhelp101):

wait never mind I stuck

OpenStudy (anonymous):

\[y=\begin{cases}x^2&\text{for }x\le0\\x+1&\text{for }x>0\end{cases}\] Take a look at the graph: